A 1500-kg truck rounds an unbanked curve on the highway at a speed of 20.0 m/s. the radius of the curvature of the curve is 80-m. calculate the coefficient of friction - is this cofficient of static or kinetic friction?
Answer #1

the coefficient must be static to avoid the car from skidding

max safe velocidad

v^2 = u r g

20^2 = u * 80* 9.8

u = 0.51

Comment before rate in case any doubt, will reply for sure.. goodluck

======

Comment before rate in case any doubt, will reply for sure.. goodluck